OpenStudy (kaliyah0524):

Algebra 1 help plz !

OpenStudy (kaliyah0524):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (agreene):

What is causing you issue? Knowing this will help me explain it to you. :)

OpenStudy (kaliyah0524):

Where to start lol

OpenStudy (agreene):

hehe okay, so there are a few ways to start. For me, as a first start, I ignore most of what a word problem says (because it doesnt matter) so with this one, the important bits are: buying flower cost $0.50 selling flower cost is 40% more how many to do you sell to get $500 So, we need something to link those together (to make x for an equation). Flowers being mentioned so much makes you want to do that... so x cost 0.50 selling x is 0.50*40% (which would be 0.50 * 0.4) how many x to get $500 how many is a good go to word for "y" So, let's make an equation x*(0.50*0.4) = 500

OpenStudy (agreene):

So, from there, we just need to simplify and then solve for x

OpenStudy (kaliyah0524):

How would i simplify that?

OpenStudy (agreene):

So, I messed it up anyway, lol

OpenStudy (agreene):

cost is 0.50 sell price is 0.50*0.4 y=500 0.50x+(0.50*0.4) = 500 the more im thinking about this the more confused im getting, partially cause the answers im getting don't really make sense.

OpenStudy (agreene):

yeah that most recent one is right.

OpenStudy (agreene):

I messed up because of the addition , its 40% more than the cost, so its 0.50*0.4 added to the price per, which is what that 0.50x+(0.50*0.4) is

OpenStudy (agreene):

so from there, you need to evaluate 0.50*4 and then solve for x

OpenStudy (agreene):

0.50*0.4 ^

OpenStudy (agreene):

actually, i think im making this harder than it is... im sorry, lol

OpenStudy (kaliyah0524):

Hold on. I gotta get a pen to work it out lol

OpenStudy (kaliyah0524):

After solving i got 2.5

OpenStudy (kaliyah0524):

So i would have the sell the flowers for $2.50 ?

OpenStudy (kaliyah0524):

@agreene